# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ Abstract
Wireless positioning technologies hold significant value for applications in autonomous driving, extended reality (XR), unmanned aerial vehicles (UAVs), and more. With the advancement of artificial intelligence (AI), leveraging AI to enhance positioning accuracy and robustness has emerged as a field full of potential. Driven by the requirements and functionalities defined in the 3rd Generation Partnership Project (3GPP) standards, AI/machine learning (ML)-based positioning is becoming a key technology to overcome the limitations of traditional methods. This paper begins with an introduction to the fundamentals of AI and wireless positioning, covering AI models, algorithms, positioning applications, emerging wireless technologies, and the basics of positioning techniques. Subsequently, focusing on standardization progress, we provide a comprehensive review of the evolution of 3GPP positioning standards, with an emphasis on the integration of AI/ML technologies in recent and upcoming releases. Based on the AI/ML-assisted positioning and direct AI/ML positioning schemes outlined in the standards, we conduct an in-depth investigation of related research. we focus on state-of-the-art (SOTA) research in AI-based line-of-sight (LOS)/non-line-of-sight (NLOS) detection, time of arrival (TOA)/time difference of arrival (TDOA) estimation, and angle estimation techniques. For Direct AI/ML Positioning, we explore SOTA advancements in fingerprint-based positioning, knowledge-assisted AI positioning, and channel charting-based positioning. Furthermore, we introduce publicly available datasets for wireless positioning and conclude by summarizing the challenges and opportunities of AI-driven wireless positioning.
@@ I Introduction
Wireless positioning, beyond its communication capabilities, represents a vital application of wireless networks, making its research and performance improvements of significant importance. In outdoor environments, cellular positioning and global navigation satellite system (GNSS) technologies play pivotal roles, providing essential support for pedestrian navigation, autonomous driving, and unmanned aerial vehicle (UAV) localization [1]. These technologies not only enhance positioning accuracy but also improve user experience and safety. Indoors, advancements in technologies such as cellular networks, WiFi, Bluetooth, and ultra-wideband (UWB) have enabled precise localization within complex environments, which is crucial for applications like shopping mall navigation and emergency evacuation [2]. As part of mobile communication infrastructure, cellular positioning is instrumental in both indoor and outdoor scenarios. However, traditional positioning methods, primarily based on geometric relationships like triangulation and trilateration, still face challenges in terms of accuracy, robustness, scalability, and adaptability to dynamic environments.
@@
In recent years, the evolution of artificial intelligence (AI) has brought transformative changes to wireless positioning [3]. AI technologies, particularly machine learning (ML) and deep learning, have significantly enhanced the accuracy and efficiency of positioning systems through their powerful data processing and pattern recognition capabilities. AI algorithms can analyze complex wireless signal characteristics, identify environmental interference factors, and predict user locations with high precision. Furthermore, AI-driven systems can adapt to changing environments by learning from historical data, offering innovative solutions to the challenges of traditional techniques. As AI continues to advance, AI-driven wireless positioning is expected to play an increasingly critical role in the future.
Nevertheless, AI-driven wireless positioning also faces several challenges. First, AI algorithms require substantial amounts of data for training, and obtaining high-quality training data in the wireless positioning domain is often difficult. Additionally, wireless signals are influenced by various factors, such as buildings and weather conditions, line-of-sight (LOS) and non-line-of-sight (NLOS) scenario, which pose challenges to the generalization capabilities of AI algorithms. Moreover, the computational complexity of AI algorithms limits their application in resource-constrained devices. Therefore, designing lightweight, efficient AI algorithms and leveraging limited data resources for effective training are crucial research areas in wireless positioning.
This survey aims to provide a comprehensive overview of AI-driven wireless positioning, encompassing the fundamentals of AI and wireless technologies, progress in standardization, state-of-the-art developments, and associated challenges and research opportunities. Although this paper mainly focuses on AI applications in cellular positioning, the survey also covers some research on AI technology on WiFi, Bluetooth, and UWB systems to present a holistic view of algorithm innovation.
@@ I-A Related Work
Wireless positioning has been a widely researched topic, with extensive investigations focusing on a variety of techniques, scenarios, and applications. Below, we provide a detailed overview of existing surveys and explain how this paper, with its focus on AI-driven wireless positioning, distinguishes itself.
@@
To address the challenges posed by wireless signal positioning, researchers have conducted numerous studies on the fundamental theories of positioning. Focusing on localization algorithms, the authors in [4] provide a comprehensive survey of time-of-arrival (TOA)-based localization algorithms. In [2], the authors summarize indoor positioning algorithms and discuss systems based on WiFi, radio frequency identification (RFID), UWB, Bluetooth, and other technologies. The survey in [5] explores fusion-based indoor positioning techniques using data from cellular networks, WiFi, global positioning systems (GPSs), inertial navigation, cameras, and more. In [6], the authors investigate device-free positioning algorithms. Regarding fingerprint-based positioning technologies, the authors in[7] and [8] conduct detailed surveys, with [9] specifically addressing methods that bypass offline fingerprint maps, and [10] delving deeper into intelligent algorithms for fingerprinting. Considering the impact of network capabilities on positioning technologies, the authors in [11] survey techniques for jammer localization in multi-hop networks, providing critical insights into handling interference and security issues in positioning systems. In [12], the authors explore advanced network localization and tracking technologies. Building on these works, the authors in [13] review the latest research in positioning technologies within wireless networks.
With advancements in technology, positioning capabilities in cellular networks have become increasingly important. In [1], the authors investigate the standardization efforts for cellular positioning from 1G to 4G and summarize the key technologies for wireless positioning in 5G networks. Following this, the paper in [14] provides an in-depth survey of positioning technologies in 5G networks, covering standardization, key elements, research trends, and performance analyses in real-world environments. Looking ahead to 6G networks, the authors in [15] summarize the latest research in positioning technologies, including novel applications, supporting technologies, system models, and critical performance indicators, while also discussing future research directions. Moreover, with advancements in communication technology, D2D networks [16], millimeter-wave (mmWave) [17], THz [18], reconfigurable intelligent surface (RIS) [19], and Ground-Air-Space Networks [20] are not only enhancing communication capabilities but also improving positioning performance. Detailed surveys on these technologically supported positioning systems have been conducted in [16, 17, 18, 19, 20].
The rapid evolution of AI has added a transformative dimension to wireless positioning. ML algorithms, in particular, have been extensively applied to extract features from wireless signals that correlate with environmental factors and user locations, thereby enabling precise and efficient positioning. Most of the above surveys incorporate discussions on ML-based positioning methodologies. Specifically, the authors in [21] provide an in-depth analysis of AI-based indoor positioning technologies. However, these studies did focus on the connection between cellular positioning and AI. Meanwhile, the swift advancements in AI have outpaced the scope of these existing reviews. This gap highlights the pressing need for an updated, comprehensive survey to support current and future research efforts.
@@ I-B Contributions
Figure 1: The overall structure of this paper.
As mentioned above, the rapid advancements in AI technology and the significant progress in AI-driven wireless positioning have motivated this comprehensive survey. This paper aims to systematically explore the fundamental principles of AI and wireless positioning technologies, ongoing standardization efforts, cutting-edge advancements, and the persistent challenges shaping future research directions. Specifically, the main contributions of this paper are as follows:
  • Foundation of AI and Wireless Positioning: This paper provides a comprehensive review of the foundational knowledge of AI and wireless positioning technologies. For AI technologies, we summarize the basic principles, introduce classical neural network models, and discuss key AI algorithms. For wireless positioning, we review various application scenarios, emerging wireless technologies, and fundamental concepts, including channel models and basic methodologies.
@@
  • 3GPP Standardization Progress: We analyze the evolution of cellular positioning within the 3rd Generation Partnership Project (3GPP) framework, summarizing its advancements from early implementations to current 5G standards. We also discuss the role of key performance indicators (KPIs) in wireless positioning and highlight the latest advancements in AI/ML-driven positioning solutions within the 3GPP standards.
@@
  • SOTA Research of AI-Driven Positioning: Based on 3GPP-defined frameworks, we summarize the state-of-the-Art (SOTA) in AI/ML-assisted positioning methods and direct AI/ML positioning methods. For AI/ML-assisted positioning methods, we focus on SOTA techniques for positioning parameter estimation, including AI-based LOS/NLOS detection, TOA/Time-difference-of-arrival (TDOA) estimation, and angle estimation algorithms. For direct AI/ML positioning methods, we classify and summarize the methods into fingerprint-based positioning, knowledge-assisted AI positioning, and channel charting based positioning, introducing the latest progress in each.
  • Datasets for Cellular Positioning: We review publicly available datasets for cellular positioning, analyzing their scenarios and unique characteristics. These datasets are essential for benchmarking AI-driven wireless positioning systems and provide critical resources for researchers seeking to validate and improve their algorithms.
@@
  • Challenges and Future Directions: We discuss the major challenges in AI-driven wireless positioning, including the scarcity of high-quality training data, the computational complexity of AI algorithms, and the model Generalization. Finally, we propose potential research directions and opportunities, providing guidance from both the perspectives of wireless technology development and AI technology advancements.
@@
Therefore, the structure of this paper is illustrated in Fig. 1. In Sec. II, we present the fundamentals of AI technology, while Sec. III focuses on the fundamentals of wireless positioning. Sec. IV summarizes the progress in 3GPP AI/ML positioning standards. The state-of-the-art advancements are reviewed in Sec. V for AI/ML-assisted positioning methods and in Sec. VI for direct AI/ML positioning methods. In Sec. VII, we investigate publicly available datasets for wireless positioning, analyzing their characteristics and application scenarios. Sec. VIII identifies the challenges and opportunities in AI-driven wireless positioning. Finally, Sec. IX concludes this paper.
II Fundamentals of AI Technology
II-A Overview of AI Models
(a) FCNN architecture.
(b) CNN architecture.
(c) LSTM architecture.
(d) Transformer architecture.
Figure 2: Architectures of the FCNN, CNN, LSTM and Transformer.
In the realm of AI, neural network architectures play a pivotal role in enabling machines to learn from data and make predictions or classifications. This subsection introduces four mainstream neural network models commonly applied in AI-driven wireless positioning systems: fully-connected neural networks (FCNNs), convolutional neural networks (CNNs), long short-term memory networks (LSTMs), and Transformers. The architecture of these models are shown in Fig. 2.
II-A1 FCNNs
FCNNs are the basic building blocks of deep learning. In an FCNN, every neuron in one layer is connected to every neuron in the subsequent layer, creating a fully connected architecture [22, 23, 3]. The network structure of a Fully Connected Neural Network (FCNN) is illustrated in Fig. 2(a). In an FCNN layer with $m$ input neurons and $n$ output neurons, each output neuron is represented as a weighted sum of all neurons from the preceding layer.
FCNNs can be employed to tackle regression, prediction, and classification tasks. When disregarding computational performance constraints, configuring deeper network models often leads to better performance. However, being a linear neural network, FCNNs inevitably encounter issues with poor accuracy when dealing with nonlinear datasets. Consequently, subsequent efforts have often focused on augmenting the nonlinear characteristics of neural networks.
II-A2 CNNs
CNNs are a type of feedforward neural network featuring convolutional operations and deep structures. They are widely used in image processing [24, 25], speech recognition [26, 27], and natural language processing [28]. As shown in Fig. 2(b), in a CNN, the convolution operation applies a set of shared weights (convolution kernels) to an input feature map to produce an output feature map.
CNNs excel at extracting local features through their convolutional layers, leveraging translation invariance to ensure efficient and accurate feature representation. Compared to FCNNs, CNNs reduce the number of parameters by sharing weights across spatial locations, making them computationally efficient for high-dimensional data [29]. However, due to their focus on local regions, CNNs inherently struggle to model global relationships, particularly at a fine-grained level. This limitation has led to the integration of complementary techniques, such as attention mechanisms or global pooling layers, to enhance their ability to capture long-range dependencies.
II-A3 LSTMs
CNNs are effective for spatial modeling but are limited in capturing temporal dependencies. To address this, LSTM networks [30, 31], a type of recurrent neural network (RNN), are designed to model sequential data by introducing gate mechanisms to regulate information flow. The schematic diagram of an LSTM is shown in Fig. 2(c). An LSTM processes input sequences using mechanisms such as the forget gate to discard irrelevant past information, the input gate to incorporate new data, and the output gate to produce meaningful hidden states.
These features enable LSTMs to retain essential long-term dependencies while adapting to new input dynamically. However, their recurrent nature and additional parameters make them computationally intensive compared to simpler feedforward networks. Beyond LSTM, other neural network models like gated recurrent unit (GRU) [32], Bidirectional RNN (BiRNN) [33], temporal convolutional network (TCN) [34], and Transformers [35] have been extensively used for time series tasks. GRU offers efficiency with fewer parameters, while BiRNN captures bidirectional dependencies. TCN achieves scalable performance by replacing recurrence with causal convolutions. Transformers leverage attention mechanisms to model long-term dependencies and are increasingly popular for time series applications.
II-A4 Transformer
Transformers are groundbreaking neural network architectures built entirely on the attention mechanism, which have revolutionized AI across multiple domains [35]. Their unified encoder-decoder structure and multimodal capabilities make them highly versatile and adaptable. As shown in Fig. 2(d), the Transformer is mainly composed of $N_{e}$ cascaded encoders and $N_{d}$ cascaded decoders. The core of the Transformer is the self-attention mechanism, which assigns dynamic weights to different parts of the input sequence, enabling the model to focus on the most relevant features [36].
Transformers streamline feature processing with a unified architecture, but their $O(n^{2})$ computational complexity for attention mechanisms imposes high resource requirements, particularly for long sequences [37]. Moreover, the reliance of Transformers on large datasets for training can limit their efficiency on smaller datasets or simpler tasks, where models like CNNs may perform better.
II-B Overview of AI Algorithms
(a) Transfer Learning.
(b) Continual Learning.
(c) Few-Shot Learning.
(d) Generative Adversarial Network.
Figure 3: Schematic diagram of Transfer learning, Continuous learning, Few-shot learning algorithms, and Generative adversarial network.
In this subsection, we will introduce the AI algorithms commonly used in wireless AI positioning, including transfer learning, continual learning, few-shot learning, and generative adversarial networks (GANs). The schematic diagram of each algorithm is shown in Fig. 3.
II-B1 Transfer Learning
Transfer learning is an AI technique that utilizes knowledge from a source domain or task to enhance learning in a target domain or task [38, 39]. This method is effective when labeled data in the target domain is limited or costly to acquire, while sufficient data or pre-trained models are available in a related source domain. As shown in Fig. 3(a), transfer learning achieves this by reusing representations from pre-trained models and adapting them to new tasks through two main steps. First, in the pre-training phase, the pre-trained model acts as a feature extractor, capturing transferable patterns from the source task. During the fine-tuning phase, the model is further trained on the target domain data, aligning it with the specific requirements of the new task by adjusting part or all of its parameters. This process enables AI models to transfer knowledge from the source domain to the target domain, significantly improving performance, especially when labeled data in the target domain is limited [39].
Transfer learning has shown great promise in wireless positioning, where collecting labeled real-world data is often challenging. For instance, knowledge from simulated datasets can be transferred to real-world scenarios, significantly reducing the need for extensive labeled data while improving system performance [40, 41, 42]. Similarly, a model trained on data from one floor of a building can be adapted to other floors, thereby lowering deployment costs. Furthermore, transfer learning can address challenges in cross-band and cross-protocol deployments, enabling models to generalize across different frequency bands or wireless communication protocols. However, the effectiveness of transfer learning in wireless positioning is closely tied to the similarity between the source and target domains. A significant discrepancy between the two domains can reduce transfer learning’s effectiveness, while a small target dataset may lead to overfitting during fine-tuning.
II-B2 Continual Learning
Continual learning [43, 44], also known as lifelong learning [45] or incremental learning [46, 47], refers to the capability of AI models to continuously learn and accumulate knowledge from past tasks, facilitating adaptation to new and emerging tasks. The schematic diagram of the continual learning algorithm is shown in Fig. 3(b). This process mimics human learning by building and retaining an evolving body of knowledge, ensuring that previously learned information is not forgotten when encountering new tasks. While continual learning shares similarities with transfer learning in leveraging prior knowledge, its objective is distinct: continual learning emphasizes long-term knowledge accumulation, whereas transfer learning focuses on optimizing a specific target task without considering long-term retention.
Continual learning is essential in wireless positioning systems, which operate in dynamic environments and constantly evolving deployment scenarios. For instance, continual learning enables positioning models to adapt to new wireless network configurations or expanded coverage areas without losing effectiveness in previously trained conditions [48]. It also addresses challenges such as adapting to evolving wireless standards, new frequency bands, changing signal propagation conditions, and hardware upgrades. Despite its potential, continual learning faces challenges, including the risk of overfitting to new tasks when data is limited and the computational overhead associated with regularization or replay-based approaches. Developing efficient and scalable continual learning methods remains a critical area for advancing wireless positioning systems.
II-B3 Few-Shot Learning
As shown in Fig. 3(c), few-shot learning aims to train effective models using a limited number of samples [49, 50]. This approach reduces the dependency on large datasets, significantly lowering the cost of data annotation and collection while maintaining competitive model performance. In scenarios where acquiring sufficient labeled data is challenging, Few-shot learning offers a promising solution by leveraging innovative learning paradigms to extract meaningful insights from sparse data. Few-shot learning models can be broadly categorized into three types: model-based, metric-based, and optimization-based approaches [49, 50].
Few-shot learning is also important for wireless positioning systems, where labeled data is often scarce, and data collection is costly. For example, it can train models for positioning in new environments or deployment configurations using minimal labeled data. However, several challenges remain. The success of Few-shot learning depends heavily on the quality and representativeness of the support set used during training. In highly diverse or complex wireless environments, identifying representative samples can be difficult. Moreover, the limited number of samples can make Few-shot learning susceptible to overfitting, especially in noisy or heterogeneous datasets. Addressing these challenges requires careful design of learning paradigms and effective strategies for generalization and robustness in wireless positioning applications.
II-B4 Generative Adversarial Networks
As shown in Fig. 3(d), GANs comprise two primary components: a Generator and a Discriminator. These two components engage in a competitive learning process where the Generator strives to produce data that closely resembles real data, while the Discriminator aims to distinguish between real and generated data. This adversarial training framework enables both models to iteratively improve their performance [51, 52].
In wireless systems, GANs are a powerful tool for data augmentation and system optimization. Their capability to generate synthetic data resembling real-world conditions is particularly beneficial in scenarios where collecting real data is expensive or logistically challenging. For example, GANs can synthesize wireless channel data under varying environmental conditions, enabling the training of robust AI models for positioning without requiring extensive real-world data collection. Moreover, GANs can assist in addressing class imbalance issues and enhancing the diversity of training datasets, thereby improving model generalization and reliability.
II-B5 Other Advanced AI Techniques
III Fundamentals of Wireless Positioning
Figure 4: AI-driven wireless positioning technologies and applications.
In this section, we first discuss the application scenarios of wireless positioning, and then list emerging wireless technologies. Finally, we introduce the basics of wireless positioning in detail, including channel models and algorithm foundations. AI-driven wireless positioning technology and applications are shown in Fig. 4.
III-A Applications of Wireless Positioning

### v2

@@ Abstract
Wireless positioning technologies hold significant value for applications in autonomous driving, extended reality (XR), unmanned aerial vehicles (UAVs), and more. With the advancement of artificial intelligence (AI), leveraging AI to enhance positioning accuracy and robustness has emerged as a field full of potential. Driven by the requirements and functionalities defined in the 3rd Generation Partnership Project (3GPP) standards, AI/machine learning (ML)-based cellular positioning is becoming a key technology to overcome the limitations of traditional methods. This paper presents a comprehensive survey of AI-driven cellular positioning. We begin by reviewing the fundamentals of wireless positioning and AI models, analyzing their respective challenges and synergies. We provide a comprehensive review of the evolution of 3GPP positioning standards, with a focus on the integration of AI/ML in current and upcoming standard releases. Guided by the 3GPP-defined taxonomy, we categorize and summarize state-of-the-art (SOTA) research into two major classes: AI/ML-assisted positioning and direct AI/ML-based positioning. The former includes line-of-sight (LOS)/non-line-of-sight (NLOS) detection, time of arrival (TOA)/time difference of arrival (TDOA) estimation, and angle prediction; the latter encompasses fingerprinting, knowledge-assisted learning, and channel charting. Furthermore, we review representative public datasets and conduct performance evaluations of AI-based positioning algorithms using these datasets. Finally, we conclude by summarizing the challenges and opportunities of AI-driven wireless positioning.
@@ I Introduction
Figure 1: Comparison of the number of wireless positioning papers and AI-driven wireless positioning papers from 2015 to 2024.
@@
Beyond communication, wireless positioning is an essential function of wireless networks. Its performance enhancement is critical for enabling a wide range of emerging applications and remains a valuable area of research. In outdoor environments, cellular positioning and global navigation satellite system (GNSS) technologies play pivotal roles, providing essential support for pedestrian navigation, autonomous driving, and unmanned aerial vehicle (UAV) localization [1]. These technologies not only enhance positioning accuracy but also improve user experience and safety. Indoors, advancements in technologies such as cellular networks, WiFi, Bluetooth, and ultra-wideband (UWB) have enabled precise localization within complex environments, which is crucial for applications like shopping mall navigation and emergency evacuation [2]. As part of the mobile communication infrastructure, cellular positioning is instrumental in both indoor and outdoor scenarios. However, traditional positioning methods, primarily based on geometric relationships like triangulation and trilateration, still face challenges in terms of accuracy, robustness, scalability, and adaptability to dynamic environments.
@@
In recent years, the evolution of artificial intelligence (AI) has brought transformative changes to wireless positioning [3]. AI technologies, particularly machine learning (ML) and deep learning, have significantly enhanced the accuracy and efficiency of positioning systems through their powerful data processing and pattern recognition capabilities [4, 5]. AI algorithms can analyze complex wireless signal characteristics, identify environmental interference factors, and predict user locations with high precision. Furthermore, AI-driven systems can adapt to changing environments by learning from historical data, offering innovative solutions to the challenges of traditional techniques. Fig. 1 illustrates the publication trends of wireless positioning-related papers over the past decade based on Google Scholar data. It can be observed that since 2017, with the rapid advancement of AI technologies, AI-driven wireless positioning has gained increasing prominence within the broader field of wireless positioning research. By 2023 and beyond, more than half of the published papers in this area have incorporated AI technologies. As AI continues to advance, AI-driven wireless positioning is expected to play an increasingly critical role in the future.
TABLE I: Comparison of Wireless Positioning Survey Papers (✓: Yes, ✗: No, ⚫: Partial)
Ref. Year Main Focus AI/ML Related 3GPP Standards Dataset Cellular Positioning
[6] 2009 TOA-based localization and NLOS mitigation techniques.
[7] 2016 WiFi fingerprint-based indoor positioning techniques.
[8] 2017 Conventional and fingerprinting-based WLAN indoor localization.
[9] 2017 Jammer localization techniques in multi-hop wireless networks.
[1] 2018 Evolution of cellular positioning technologies from 1G to 5G.
[10] 2018 Network localization, tracking, and navigation techniques.
[11] 2019 Indoor localization approaches without offline fingerprint maps.
[2] 2019 Indoor localization techniques in WiFi, UWB, and RFID systems.
[12] 2019 Device-free localization for smart environment applications.
[13] 2020 Fusion-based localization systems for indoor positioning.
[14] 2020 ML-enhanced fingerprint-based indoor localization.
[15] 2020 ML technology for wireless positioning.
[16] 2021 D2D-based cooperative positioning techniques.
[17] 2022 THz localization techniques for future 6G networks.
[18] 2022 mmWave localization and sensing technologies.
[19] 2022 RIS-assisted localization methods for IoT applications.
[20] 2023 High-accuracy localization technologies for 6G networks.
[21] 2024 Comprehensive 5G positioning systems and standards.
[22] 2024 Localization techniques for integrated ground-air-space networks.
[23] 2024 Comprehensive survey of indoor wireless positioning technologies.
    Our Work AI-driven positioning techniques for cellular networks, including fundamentals, 3GPP standards, SOTA, datasets, and challenges.
@@ I-A Related Work
Wireless positioning is a widely researched topic, with extensive investigations focusing on a variety of techniques, scenarios, and applications. Below, we provide a detailed overview of existing surveys and explain how this paper, with its focus on AI-driven wireless positioning, distinguishes itself.
@@
To address the challenges posed by wireless positioning, researchers have conducted numerous studies on the fundamental theories of positioning. Focusing on localization algorithms, the authors in [6] provide a comprehensive survey of time-of-arrival (TOA)-based localization algorithms. In [2], the authors summarize indoor positioning algorithms and discuss systems based on WiFi, radio frequency identification (RFID), UWB, Bluetooth, and other technologies. The survey in [13] explores fusion-based indoor positioning techniques using data from cellular networks, WiFi, global positioning systems (GPSs), inertial navigation, cameras, and more. In [12], the authors investigate device-free positioning algorithms. Regarding fingerprint-based positioning technologies, the authors in[7] and [8] conduct detailed surveys, with [11] specifically addressing methods that bypass offline fingerprint maps, and [14] delving deeper into intelligent algorithms for fingerprinting. Considering the impact of network capabilities on positioning technologies, the authors in [9] survey techniques for jammer localization in multi-hop networks, providing critical insights into handling interference and security issues in positioning systems. In [10], the authors explore advanced network localization and tracking technologies. Building on these works, the authors in [23] review the latest research in positioning technologies within wireless networks. With advancements in technology, positioning capabilities in cellular networks have become increasingly important. In [1], the authors investigate the standardization efforts for cellular positioning from 1G to 4G and summarize the key technologies for wireless positioning in 5G networks. Following this, the paper in [21] provides an in-depth survey of positioning technologies in 5G networks, covering standardization, key elements, research trends, and performance analyses in real-world environments. Looking ahead to 6G networks, the authors in [20] summarize the latest research in positioning technologies, including novel applications, supporting technologies, system models, critical performance indicators, and future research directions. Moreover, with advancements in communication technology, D2D networks [16], millimeter-wave (mmWave) [18], THz [17], reconfigurable intelligent surface (RIS) [19], and Ground-Air-Space Networks [22] are not only enhancing communication capabilities but also improving positioning performance.
The summary and comparison of existing survey papers are provided in Table. I. Some prior surveys have mentioned AI-related techniques for wireless positioning [8, 10, 2, 12, 13, 17, 18, 20, 21, 22, 23]. However, such discussions are typically brief and lack a systematic review of the technical landscape and evolution of AI/ML models in this domain. Among existing works, the survey papers in [14, 15] present relatively comprehensive reviews of AI/ML applications in wireless positioning. However, there are substantial differences between their works and ours. The survey in [14] primarily focuses on traditional ML techniques for positioning, while our work focuses more on advanced AI techniques (e.g., deep learning methods). The survey in [15] focuses on different AI technologies and discusses the roles of various AI technologies in positioning tasks. In contrast, our work focuses specifically on different types of positioning algorithms (such as AI/ML-assisted positioning and direct AI/ML positioning methods), and discusses how AI can enhance each of these categories. For wireless positioning standards, numerous studies have reviewed the progress of 3GPP standards [1, 16, 20, 21]. Nonetheless, the investigation into the integration of AI techniques within the 3GPP standardization framework remains insufficient. Our survey includes an in-depth analysis of AI-enabled positioning within the 3GPP framework. Regarding datasets, the survey in [15] broadly covers datasets for WiFi, Bluetooth, and other wireless technologies, with minimal attention given to cellular positioning datasets. Conversely, we place significant emphasis on datasets pertinent to cellular positioning. Similar to works such as [1, 16, 17, 18, 20, 21], our survey focuses on the domain of cellular positioning. Considering the rapid advancements in AI for wireless positioning over the past five years (as shown in Fig. 1), we concentrate on reviewing AI-driven positioning solutions in cellular networks, aiming to provide targeted, up-to-date, and comprehensive insights into this fast-evolving field. Specifically, our goal is to comprehensively review cutting-edge AI-driven wireless positioning techniques for cellular networks, including their fundamentals, 3GPP standards, state-of-the-art (SOTA) algorithms, datasets, and open challenges.
Figure 2: The overall structure of this paper.
@@ I-B Contributions
As mentioned above, the rapid advancements in AI technology and the significant progress in AI-driven wireless positioning have motivated this comprehensive survey. The core contribution of our survey is to provide a focused and structured review of AI-driven cellular positioning, covering its fundamentals, standards, SOTA developments, datasets, and key challenges. The specific contributions of this paper are as follows:
  • Foundation of AI and Wireless Positioning: We present a comprehensive review of the foundational knowledge of both wireless positioning and AI technologies. For wireless positioning, we review key application scenarios, wireless channel models, and positioning algorithms. We further analyze the limitations of model-based positioning approaches and motivate the shift toward AI-driven solutions. On the AI side, we review representative neural network models and algorithms, and analyze their respective strengths, limitations, and suitability for different positioning tasks.
@@
  • 3GPP Standardization Progress: We analyze the evolution of cellular positioning within the 3rd Generation Partnership Project (3GPP) framework, summarizing its advancements from early implementations to current 5G standards. We discuss the role of key performance indicators (KPIs) in wireless positioning and highlight the latest advancements in AI/ML-driven positioning solutions within the 3GPP standards. Moreover, we discuss potential future directions for AI-driven positioning in the upcoming 6G and beyond standards.
@@
  • SOTA Research of AI-Driven Positioning [^1] While our primary focus is on cellular-based positioning, we also selectively include non-cellular techniques in the SOTA review, to ensure completeness and to highlight transferable methodologies across different wireless technologies.: Based on 3GPP-defined frameworks, we summarize the SOTA in AI/ML-assisted positioning methods and direct AI/ML positioning methods. For AI/ML-assisted positioning methods, we focus on SOTA techniques for positioning parameter estimation, including AI-based line-of-sight (LOS)/non-line-of-sight (NLOS) scenarios detection, TOA/Time-difference-of-arrival (TDOA) estimation, and angle estimation algorithms. For direct AI/ML positioning methods, we classify and summarize the methods into fingerprint-based positioning, knowledge-assisted AI positioning, and channel charting based positioning, introducing the latest progress in each.
  • Datasets for Cellular Positioning: We review publicly available datasets relevant to AI-based cellular positioning, summarizing their use cases, data characteristics, and limitations. In particular, we analyze two representative datasets (measurement-based MaMIMO and synthetic-based DeepMIMO) as case studies to evaluate AI models’ performance across different deployment scenarios. To facilitate reproducibility and accelerate research in this area, the corresponding implementation code has been released as open source on GitHub  [^2] The code is available at: https://github.com/guangjinpan/AI-Driven-Localization.. We also identify current gaps in dataset availability and propose future directions for constructing more diverse and scalable datasets.
@@
  • Challenges and Future Directions: We discuss the major challenges in AI-driven wireless positioning, including data collection, accuracy in complex environments, and model generalization. We also present potential solutions to these challenges. Finally, we propose possible research directions and opportunities, providing guidance from both the perspectives of wireless technology development and AI technology advancements.
@@
Therefore, the structure of this paper is illustrated in Fig. 2. In Sec. II, we present the fundamentals of wireless positioning, while Sec. III focuses on the fundamentals of AI technology. Sec. IV summarizes the progress in 3GPP AI/ML positioning standards. The SOTA advancements are reviewed in Sec. V for AI/ML-assisted positioning methods and in Sec. VI for direct AI/ML positioning methods. In Sec. VII, we investigate publicly available datasets for wireless positioning, analyzing their characteristics and application scenarios. Sec. VIII identifies the challenges and opportunities in AI-driven wireless positioning. Finally, Sec. IX concludes this paper.
II Fundamentals of Wireless Positioning
Wireless positioning is a key enabler for emerging applications. However, achieving high-precision positioning remains challenging due to the complex and dynamic nature of wireless propagation environments. In this section, we introduce the fundamentals of wireless positioning, focusing on real-world requirements and challenges and the motivation for AI integration. Specifically, we first review critical application scenarios that demand accurate positioning. Then, we present basic positioning principles, covering models of key measurements and positioning methods (including ranging-based, angle-based, fingerprint-based, and channel charting-based solutions). Finally, we deeply analyze the challenges faced by model-based approaches and discuss how AI-driven techniques can address them, providing a comparative analysis to motivate the integration of AI into positioning systems.
II-A Applications of Wireless Positioning